Jhansy Garza's regular hourly wage rate is $15.00, and she receives a wage of 1½ times the regular hourly rate for work in excess of 40 hours. During a March weekly pay period, Jhansy worked 43 hours. Her gross earnings prior to the current week were $6,500. Jhansy is married and claims three withholding allowances. Her only voluntary deduction is for group hospitalization insurance at $40 per week.

MARRIED Persons ― WEEKLY Payroll Period
(For Wages Paid through December 2017)

If the wages are ―

And the number of withholding allowances claimed is ―

At least

But less
than

0 1 2 3 4 5 6 7 8 9 10
The amount of income tax to be withheld is ―
500

510

34 27 19 11 4 0 0 0 0 0

0

510

520

35 28 20 12 5 0 0 0 0 0

0

520

530

37 29 21 13 6 0 0 0 0 0

0

530

540

38 30 22 14 7 0 0 0 0 0

0

540

550

40 31 23 15 8 0 0 0 0 0

0

550

560

41 32 24 16 9 1 0 0 0 0

0

560

570

43 33 25 17 10 2 0 0 0 0

0

570

580

44 34 26 18 11 3 0 0 0 0

0

580

590

46 35 27 19 12 4 0 0 0 0

0

590

600

47 36 28 20 13 5 0 0 0 0

0

600

610

49 38 29 21 14 6 0 0 0 0

0

610

620

50 39 30 22 15 7 0 0 0 0

0

620

630

52 41 31 23 16 8 1 0 0 0

0

630

640

53 42 32 24 17 9 2 0 0 0

0

640

650

55 44 33 25 18 10 3 0 0 0

0

650

660

56 45 34 26 19 11 4 0 0 0

0

660

670

58 47 35 27 20 12 5 0 0 0

0

670

680

59 48 37 28 21 13 6 0 0 0

0

680

690

61 50 38 29 22 14 7 0 0 0

0

690

700

62 51 40 30 23 15 8 0 0 0

0

a. Compute the following amounts for Jhansy’s wages for the current week. (Round FICA taxes, State income taxes and Net pay answers to 2 decimal places, e.g. 52.55.)

(1) Gross earnings. $enter a dollar amount
(2) FICA taxes. (Assume a 7.65% rate on maximum of $127,200.) $enter a dollar amount rounded to 2 decimal places
(3) Federal income taxes withheld. $enter a dollar amount
(4) State income taxes withheld. (Assume a 2.0% rate.) $enter a dollar amount rounded to 2 decimal places
(5) Net pay.

b. Record Jhansy’s pay. (Credit account titles are automatically indented when amount is entered. Do not indent manually. Round answers to 2 decimal places, e.g. 52.75.)
Answer #1
Normal Pay (40*15) $      600.00
Overtime Pay (Overtime hours = 43 - 40 =3) (Overtime rate = 15*1.5 = 22.5) (3*22.5) $         67.50
Total Gross pay (Rounded to Dollar Amount) $            668
Total Gross Pay till the Date = 6500+667.50 = 7167.5 which is less than maximum limit of 127,200)
FICA taxes (668*7.65%) $         51.10
Federal income taxes withheld (where claims three withholding allowances and 667.50 gross pay) $               27
State income taxes withheld (667.50*2%) $         13.35
Total Gross pay $      668.00
Less: Deductions
Group hospitalization insurance $                 40
FICA taxes $           51.10
Federal income taxes withheld $                 27
State income taxes withheld $           13.35 $      131.45
Net pay $      536.55
Date Accounts title and explanations Debit Credit
Salaries and Wages Expense $               668
           FICA taxes payable $         51.10
           Federal Income taxes payable $               27
           State Income Tax payable $         13.35
           Group Insurance Payable $               40
           Salaries and Wages Payable $      536.55
To record the Jhansy’s pay.